# V2X-Radar: A Multi-modal Dataset with 4D Radar for Cooperative Perception

### v1

@@ Abstract
Modern autonomous vehicle perception systems often struggle with occlusions and limited perception range. Previous studies have demonstrated the effectiveness of cooperative perception in extending the perception range and overcoming occlusions, thereby improving the safety of autonomous driving. In recent years, a series of cooperative perception datasets have emerged. However, these datasets only focus on camera and LiDAR, overlooking 4D Radar, a sensor employed in single-vehicle autonomous driving for robust perception in adverse weather conditions. In this paper, to bridge the gap of missing 4D Radar datasets in cooperative perception, we present V2X-Radar, the first large real-world multi-modal dataset featuring 4D Radar. Our V2X-Radar dataset is collected using a connected vehicle platform and an intelligent roadside unit equipped with 4D Radar, LiDAR, and multi-view cameras. The collected data includes sunny and rainy weather conditions, spanning daytime, dusk, and nighttime, as well as typical challenging scenarios. The dataset comprises 20K LiDAR frames, 40K camera images, and 20K 4D Radar data, with 350K annotated bounding boxes across five categories. To facilitate diverse research domains, we establish V2X-Radar-C for cooperative perception, V2X-Radar-I for roadside perception, and V2X-Radar-V for single-vehicle perception. We further provide comprehensive benchmarks of recent perception algorithms on the above three sub-datasets. The dataset and benchmark codebase will be available at: http://openmpd.com/column/V2X-Radar.
Figure 1: A data frame sampled from V2X-Radar dataset. Each sample consists of data from three distinct sensors: (1) Dense point clouds (gray points) from roadside and vehicle-side LiDAR. (2) Sparse point cloud (green / blue points) with doppler information from roadside and vehicle-side 4D Radar. (3) RGB images (top row) from vehicle‘s camera and roadside multi-view cameras. All sensors have been synchronized in both time and spatial dimensions. The data frame is further annotated manually with 3D box across five categories, each category being represented by a unique color.
@@ 1 Introduction
Perception is critical in autonomous driving, and while a lot of single-vehicle perception methods [25, 24, 12, 14, 22, 48, 39, 26, 20] have emerged, it still faces substantial safety challenges arising from occlusions and limited perception ranges. These issues arise because vehicles can only see their surroundings from a single view, leading to an incomplete scenario perception. This limitation hinders autonomous vehicles from achieving safe navigation and making optimal decisions. To address this limitation, recent works have delved into the vehicle-to-everything (V2X) cooperative perception, where the ego vehicle utilizes the information from other agents through wireless communication to extend perception range and overcome occlusions.
@@
Recently, cooperative perception has gained increasing attention, and several pioneering datasets have been released to support this research. For instance, datasets like OpenV2V [33], V2X-Sim [10], and V2XSet [32] are generated through simulations using CARLA [3] and SUMO [8]. Conversely, datasets such as DAIR-V2X [42], V2X-Seq [43], V2V4Real [35], and V2X-Real [28] are derived from real-world scenarios. However, a shared limitation of these datasets is their exclusive focus on camera and LiDAR sensors, overlooking the potential of 4D Radar. This sensor is considered helpful to robust perception for its exceptional adaptability to adverse weathers, which has been validated in the single-vehicle autonomous driving datasets such as K-Radar [17] and Dual-Radar [46].
@@
In order to fill the gap of missing 4D Radar dataset in the field of cooperative perception and thus facilitate related studies, we present V2X-Radar, the first large real-world multi-modal dataset with 4D Radar. This dataset comprises a wide variety of scenarios, encompassing various weather conditions such as sunshine, rain and snow, spanning different times of the day including daytime, dusk, and nighttime. The data is collected using a connected vehicle platform and an intelligent roadside unit, both equipped with 4D Radar, LiDAR, and multi-view cameras (Fig. 1). From over 15 hours of collected driving logs, we carefully selected 50 representative scenarios for the final dataset. It comprises 20K LiDAR frames, 40K camera images, and 20K 4D Radar data, featuring 350K annotated bounding boxes across 5 object categories. To support a variety of research areas, the V2X-Radar is further divided into three specialized sub-datasets: V2X-Radar-C for cooperative perception, V2X-Radar-I for roadside perception, and V2X-Radar-V for single-vehicle perception. Additionally, we provide comprehensive benchmarks of recent perception algorithms across these three sub-datasets. Compared to the existing real-world cooperative datasets, our V2X-Radar shows two strengths: (1) More modalities: The proposed V2X-Radar dataset includes three types of sensors: LiDAR, Camera, and 4D Radar, enabling further exploration into 4D Radar-related cooperative perception research. (2) Diverse scenarios: Our data collection encompasses various weather conditions and times of day, while also focuses on challenging intersections for single-vehicle autonomous driving. These intersections have obstructed blind spots that significantly impact vehicle safety, providing diverse corner cases for cooperative perception research.
@@
  • We build V2X-Radar, the first large real-world multi-modal dataset featuring 4D Radar tailored for cooperative perception. All the frames are captured through diverse real-world scenarios using multi-modal sensors.
  • We provide 20K LiDAR frames, 40K multi-view camera images, and 20K 4D Radar data, alongside 350K annotated bounding boxes spanning five object categories.
  • Comprehensive benchmarks of recent perception algorithms are conducted across cooperative perception on V2X-Radar-C, roadside perception on V2X-Radar-I, and vehicle-side perception on V2X-Radar-V. We will release all the V2XS-Radar dataset and benchmark codebase.
Dataset Year Location Sensor V2X Multi-view Cameras Real Data 4D Radar
KITTI [5] 2012 Germany L&C
Waymo [23] 2019 USA L&C
nuScenes [1] 2019 USA L&C&R
OpenV2V [33] 2022 Carla L&C V2V
V2X-Sim [10] 2022 Carla L&C V2V&I
V2X-Set [32] 2022 Carla L&C V2V&I
DAIR-V2X [42] 2022 China L&C V2I
V2X-Seq [43] 2023 China L&C V2I
V2V4Real [35] 2023 USA L&C V2V
V2X-Real [28] 2024 USA L&C V2V&I
TUMTraf-V2X [50] 2024 Germany L&C V2I
[…]
@@ 2.1 Autonomous Driving Datasets
Public datasets [1, 23, 33, 32, 10, 43, 42, 35, 28, 17, 46, 47] have greatly accelerated the progress of autonomous driving in recent years. See Tab. 1 for a summary of autonomous driving datasets. Single-vehicle datasets, such as KITTI [5], NuScenes [1], and Waymo [23], have significantly promoted the development of single-vehicle perception. For the sake of cooperative perception datasets, OPV2V [33] is the first dataset in this field. It collects data through CARLA [3] and OpenCDA [30, 34] co-simulation. V2XSet [32] and V2X-Sim [10] explore V2X perception using synthesized data from the CARLA simulator [3]. In contrast to simulated datasets, DAIR-V2X introduces the first real-world dataset for cooperative detection. V2X-Seq [43] extends sequences from DAIR-V2X [42] with track IDs, creating a sequential perception and trajectory forecasting dataset. V2V4Real [35] presents the initial real-world V2V dataset collected by two connected vehicles.V2X-Real [28] is a large-scale multi-modal multi-view dataset designed for V2X research, gathered using two connected automated vehicles and two intelligent roadside units. A common limitation among the above cooperative perception datasets is their emphasis solely on camera and LiDAR sensors, overlooking the capabilities of 4D Radar. 4D Radar is known for its superior adaptability to adverse weather conditions, which has been proven by single-vehicle autonomous driving datasets like K-Radar [17] and Dual-Radar [46]. Therefore, it is crucial to introduce a real-world cooperative dataset incorporating 4D Radar to fill this gap and further facilitate 4D Radar-related cooperative perception research.
@@ 2.2 Cooperative Perception
Cooperative perception aims to extend the perception range and overcome occlusions in single-vehicle perception by leveraging the shared information among connected agents. It can be classified into three main types based on fusion strategies.: (1) Early Fusion, which involves transmitting raw sensor data for the ego-vehicle to aggregate and predict objects. (2) Late Fusion that integrates detection results for a consistent prediction. (3) Intermediate Fusion that shares and fuses intermediate high-level features.
Recent state-of-the-art methods [27, 2, 31, 32, 6, 16, 49, 27, 4, 45, 44] typically fellow the intermediate fusion, which achieves the best trade-off between accuracy and bandwidth requirements. For instance, V2VNet [27] utilized graph neural networks for integrating intermediate features. F-Cooper [2] applied max-pooling fusion to merge shared voxel features. CoBEVT [31] introduced local-global sparse attention for improved cooperative BEV map segmentation. V2X-ViT [32] presented a unified vision transformer for multi-agent and multi-scale perception. HM-ViT [6] introduced a 3D heterogeneous graph transformer for camera and LiDAR fusion. HEAL [16] proposed an extensible collaborative perception framework to integrate new heterogeneous agents with minimal integration cost and performance decline.
@@ 3 V2X-Radar Dataset
To facilitate the cooperative perception research with 4D Radar, we present V2X-Radar, the first large real-world multi-modal dataset featuring 4D Radar. In this section, we first detail the data acquisition in Sec. 3.1, then present the data annotation in Sec. 3.2, and finally delve into the diverse data distribution and dataset analysis in Sec. 3.3.
@@ 3.1 Data Acquisition
Sensor Setup. The dataset was collected using a connected vehicle-side platform (Fig. 2(a)) and an intelligent roadside unit (Fig. 2(b)). Both are equipped with sensors such as 4D Radar, LiDAR, and multi-view cameras. Additionally, the GPS/IMU system is employed to achieve high-precision localization, which serves for the initial point cloud registration between vehicle-side and roadside platforms. C-V2X unit is also integrated for wireless data transmission. The sensor layout configuration can be found in Fig. 2, and detailed specifications are listed in Tab. 2.
Figure 2: The sensor configuration on the connected vehicle-side platform and the intelligent roadside unit. a) The vehicle-side platform. b) The intelligent roadside unit or infrastructure unit. Both of them are equipped with multi-modal sensors including camera, LiDAR and 4D Radar, along with a C-V2X unit and a GPS/IMU system.
@@
Synchronization. For cooperative perception datasets, it is crucial to synchronize sensors on both vehicle-side and roadside platforms using a uniform timestamp standard. To ensure consistency, all computer clocks are first aligned with GPS time. Hardware-triggered synchronization of LiDAR, cameras, and 4D Radar is then implemented using the Precision Time Protocol (PTP) and Pulse Per Second (PPS) signals. Subsequently, the closest LiDAR frames from the vehicle-side and roadside platforms are matched, and camera and 4D Radar data are aligned with each corresponding LiDAR frame to create unified multi-modal data frames. Finally, the time difference between sensors across the two platforms is maintained below 50 milliseconds for each sample.
Sensor Calibration and Registration. Through the sensor calibration process, we achieve spatial synchronization of the camera, LiDAR, and 4D Radar. The intrinsic parameters of the camera are calibrated using a checkerboard pattern. In contrast, the LiDAR is calibrated relative to the camera by utilizing 100 point pairs extracted from the point cloud and the corresponding camera image. The extrinsic parameters are derived by minimizing the reprojection errors between the 2D-3D point correspondences. The calibration of the LiDAR with the 4D Radar is performed by selecting 100 high-intensity point pairs located on corner reflectors. The results of this calibration are visually represented in Fig. 3. Vehicle-side LiDAR alignment with roadside LiDAR is achieved through point cloud registration, initially computed using RTK localization and subsequently refined via CBM [21] and manual adjustments. The visualization of the point cloud registration is shown in Fig. 4.
Figure 3: Visualization of calibration results. a) The calibration results between the camera and LiDAR. b) The calibration results between 4D Radar and LiDAR / Camera. The LiDAR points are projected onto the camera plane using camera intrinsics and camera-LiDAR extrinsic. Similarly, the 4D Radar points are transferred onto the LiDAR coordinate system using the 4D Radar-LiDAR extrinsic. Additionally, these 4D Radar points are also mapped onto the camera plane by leveraging the camera’s intrinsic parameters, along with the extrinsic parameters of camera-LiDAR and 4D Radar-LiDAR.
@@
Data Collection. We collected 15 hours of driving data with 540K frames, capturing various weather conditions such as sunny and rainy, spanning daytime, dusk, and nighttime, as well as typical challenging intersection scenarios. Further details can be found in the supplementary materials. From this, we manually select the most representative 40 sequences to form V2X-Radar-C; each sequence is 10-25 seconds long with a 10Hz capturing frequency. On this basis, we additionally sample 10 vehicle-only sequences to create V2X-Radar-V and sample an additional 10 infrastructure-only sequences to form V2X-Radar-I. In comparison to the single-view data in V2X-Radar-C, both V2X-Radar-V and V2X-Radar-I encompass more diverse scenes. Consequently, the above three sub-datasets collectively encompassed 20K LiDAR frames, 40K camera images, and 20K 4D Radar data.
@@
Coordinate System. Our dataset comprises four types of coordinate systems: (1) LiDAR coordinate system, where the X, Y, and Z axes align with the front, left, and upward directions of the LiDAR. (2) Camera coordinate systems, where the z-axis denotes depth. (3) 4D Radar coordinate system, where the X, Y, and Z axes are oriented towards the right, front, and upward directions. (4) Global coordinate system, which maintains consistency with the LiDAR frame on the roadside platform.
Figure 5: Data analysis of our V2X-Radar dataset. a) Objects distribution during day and night conditions. b) Average and maximum number of LiDAR points within the 3D bounding box for each category. c) Average and maximum number of 4D Radar points within the 3D bounding box for each category. d) Number of annotations per collaborative sample. The vertical axes of sub-plots (a)-(c) employ a log scale, while the vertical axis of sub-plot (d) utilizes a standard scale.
3D Bounding Boxes Annotation.
Data annotation includes annotations for vehicles, roadsides, and collaborative efforts. Vehicle and roadside annotations are manually created within their specific LiDAR coordinate systems. For the collaborative annotation, we first align the vehicle-side and roadside annotations within a unified roadside LiDAR coordinate system, then match them using the Intersection over Union (IoU) metric to eliminate duplicates. The annotation process involves five object categories: pedestrian, cyclist, car, bus, and truck. Each object annotation includes attributes such as occlusion and truncation, as well as the geometric properties of 3D bounding boxes, denoted as $(x,y,z,w,h,l,\theta)$, where $(l.w.h)$ represent the object’s dimensions, $(x,y,z)$ indicate its location, and $\theta$ denotes its orientation.

### v2

@@ Abstract
Modern autonomous vehicle perception systems often struggle with occlusions and limited perception range. Previous studies have demonstrated the effectiveness of cooperative perception in extending the perception range and overcoming occlusions, thereby enhancing the safety of autonomous driving. In recent years, a series of cooperative perception datasets have emerged; however, these datasets primarily focus on cameras and LiDAR, neglecting 4D Radar—a sensor used in single-vehicle autonomous driving to provide robust perception in adverse weather conditions. In this paper, to bridge the gap created by the absence of 4D Radar datasets in cooperative perception, we present V2X-Radar, the first large-scale, real-world multi-modal dataset featuring 4D Radar. V2X-Radar dataset is collected using a connected vehicle platform and an intelligent roadside unit equipped with 4D Radar, LiDAR, and multi-view cameras. The collected data encompasses sunny and rainy weather conditions, spanning daytime, dusk, and nighttime, as well as various typical challenging scenarios. The dataset consists of 20K LiDAR frames, 40K camera images, and 20K 4D Radar data, including 350K annotated boxes across five categories. To support various research domains, we have established V2X-Radar-C for cooperative perception, V2X-Radar-I for roadside perception, and V2X-Radar-V for single-vehicle perception. Furthermore, we provide comprehensive benchmarks across these three sub-datasets. We will release all datasets [^1] http://openmpd.com/column/V2X-Radar and benchmark codebase [^2] https://github.com/yanglei18/V2X-Radar.
Figure 1: A data frame sampled from the V2X-Radar dataset. Each sample includes data from three sensors: (1) dense point clouds (gray points) from roadside and vehicle-side LiDAR; (2) sparse point clouds (green and blue points) with Doppler information from the roadside and vehicle-mounted 4D Radar; (3) RGB images (top row) from the vehicle’s camera and multi-view roadside cameras. All sensors are temporally and spatially synchronized. Each data frame is manually annotated with 3D boxes across five categories.
Dataset Reference Location Sensor V2X Real 4D Radar Adverse Weather Day&Night #LiDAR #Images #4D Radar #3D Box
OpenV2V [37] ICRA’22 Carla L&C V2V 11K 44K 0 233K
V2X-Set [36] ECCV’22 Carla L&C V2V&I 11K 44K 0 233K
DAIR-V2X [48] CVPR’22 China L&C V2I 39K 39K 0 464K
V2X-Seq [49] CVPR’23 China L&C V2I 39K 39K 0 464K
V2V4Real [39] CVPR’23 USA L&C V2V 20K 40K 0 240K
RCooper [7] CVPR’24 China L&C I2I 30K 50K 0 310K
TUMTraf-V2X[55] CVPR’24 DE L&C V2I 2K 5K 0 29.38K
[…]
@@ 1 Introduction
Perception is critical in autonomous driving. While many single-vehicle perception methods [29, 28, 16, 26, 54, 43, 23, 23, 22, 25, 14] have emerged, substantial safety challenges persist due to occlusions and limited perception ranges. These issues occur because vehicles can only view their surroundings from a single perspective, resulting in an incomplete understanding of the scenario. This limitation hinders autonomous vehicles from achieving safe navigation and making optimal decisions. To address this challenge, recent studies have explored cooperative perception, where ego vehicles extend their perception range and overcome occlusions with assistance from other vehicles or roadside perception [42, 41, 44, 45, 31].
@@
Recently, cooperative perception has attracted increasing attention, and several pioneering datasets have been released to bolster this research. For instance, datasets like OpenV2V [37], V2X-Sim [12], and V2XSet [36] are generated through simulations using CARLA [3] and SUMO [10]. In contrast, datasets such as DAIR-V2X [48], V2X-Seq [49], V2V4Real [39], and V2X-Real [32] are derived from real-world scenarios. However, a common limitation of these datasets is their exclusive focus on camera and LiDAR sensors, neglecting the potential of 4D Radar. This sensor is considered advantageous for robust perception due to its exceptional adaptability to adverse weather conditions, as evidenced in single-vehicle autonomous driving datasets such as K-Radar [19] and Dual-Radar [52].
@@
To address the gap in the 4D Radar dataset for cooperative perception and to facilitate related studies, we present V2X-Radar, the first large-scale real-world multi-modal dataset featuring 4D Radar. This dataset includes a diverse range of scenarios, covering various weather conditions such as sunny, rainy, and snowy environments, and spans different times of the day, including daytime, dusk, and nighttime. The data was collected using a connected vehicle platform alongside an intelligent roadside unit, both equipped with 4D Radar, LiDAR, and multi-view cameras (Fig. 1). From over 15 hours of driving logs, we meticulously selected 50 representative scenarios for the final dataset. It comprises 20K LiDAR frames, 40K camera images, and 20K 4D Radar data, featuring 350K annotated bounding boxes across five object categories. To support a variety of research areas, V2X-Radar is further divided into three specialised sub-datasets: V2X-Radar-C for cooperative perception, V2X-Radar-I for roadside perception, and V2X-Radar-V for single-vehicle perception. Additionally, we provide comprehensive benchmarks of recent perception algorithms across these three sub-datasets. In comparison to existing real-world cooperative datasets, our V2X-Radar demonstrates two key strengths: (1) More modalities: The proposed V2X-Radar dataset includes three types of sensors: LiDAR, Camera, and 4D Radar, enabling further exploration into 4D Radar-related cooperative perception research. (2) Diverse scenarios: Our data collection covers diverse weather conditions and times of day, focusing on complex intersections that pose challenges for single-vehicle autonomous driving. These scenarios feature obstructed blind spots that impact vehicle safety, providing varied corner cases for cooperative perception research.
@@
  • We have developed V2X-Radar, the first extensive real-world multi-modal dataset featuring 4D Radar designed for cooperative perception. All frames are captured through a variety of real-world scenarios using multi-modal sensors.
  • We offer 20K LiDAR frames, 40K multi-view camera images, and 20K 4D Radar data, accompanied by 350K annotated bounding boxes across five object categories.
  • Comprehensive benchmarks of recent perception algorithms are conducted across cooperative perception on V2X-Radar-C, roadside perception on V2X-Radar-I, and vehicle-side perception on V2X-Radar-V. We will release the entire V2XS-Radar dataset and benchmark codebase.
@@ 2.1 Autonomous Driving Datasets
Public datasets [1, 27, 37, 36, 12, 49, 48, 39, 32, 19, 52, 53] have significantly accelerated the progress of autonomous driving in recent years. Single-vehicle datasets, such as KITTI [6], NuScenes [1], and Waymo [27], have notably furthered the development of single-vehicle perception. In terms of cooperative perception datasets, OPV2V [37] is the first dataset in this area, gathering data through co-simulation using CARLA [3] and OpenCDA [34, 38]. V2XSet [36] and V2X-Sim [12] explore V2X perception by employing synthesized data from the CARLA simulator [3]. Unlike simulated datasets, DAIR-V2X introduces the first real-world dataset for cooperative detection. V2X-Seq [49] extends sequences from DAIR-V2X [48] with track IDs, establishing a sequential perception and trajectory forecasting dataset. V2V4Real [39] presents the initial real-world V2V dataset gathered from two connected vehicles. V2X-Real [32] is a large-scale multi-modal multi-view dataset intended for V2X research, compiled using two connected automated vehicles and two intelligent roadside units. TUMTraf-V2X [55] is a multimodal, multi-view V2X cooperative perception dataset designed for 3D object detection and tracking in traffic scenarios. A common limitation among the cooperative perception datasets mentioned above is their narrow focus on camera and LiDAR sensors, overlooking the benefits of 4D Radar. 4D Radar is recognized for its superior adaptability to adverse weather conditions, as demonstrated by single-vehicle autonomous driving datasets such as K-Radar [19] and Dual-Radar [52]. Consequently, It’s essential to create a comprehensive multimodal dataset that integrates 4D Radar to enhance cooperative perception research.
@@ 2.2 Cooperative Perception
Cooperative perception aims to extend the range of perception and overcome occlusions in single-vehicle perception by leveraging shared information among connected agents. It can be classified into three main types based on fusion strategies: (1) Early Fusion, which involves transmitting raw sensor data for the ego vehicle to aggregate and predict objects; (2) Late Fusion, which integrates detection results for a consistent prediction; and (3) Intermediate Fusion, which shares and fuses intermediate high-level features. Recent state-of-the-art methods [30, 2, 35, 36, 8, 18, 30, 4, 50, 51, 5] typically follow intermediate fusion, achieving the best trade-off between accuracy and bandwidth requirements. For instance, V2VNet [30] uses a graph neural network to refine features and perform joint perception and prediction iteratively. F-Cooper [2] introduces a pooling mechanism for identifying salient features. CoBEVT [35] introduced local-global sparse attention for improved cooperative BEV map segmentation. V2X-ViT [36] presented a unified vision transformer for multi-agent and multi-scale perception. HM-ViT [8] introduced a 3D heterogeneous graph transformer for camera and LiDAR fusion. FFNet [50] is a novel flow-based feature fusion framework that addresses temporal asynchrony in cooperative 3D object detection through feature flow prediction. HEAL [18] proposed an extensible collaborative perception framework to integrate new heterogeneous agents with minimal integration cost and performance decline.
@@ 3 V2X-Radar Dataset
To facilitate cooperative perception research using 4D Radar, we introduce V2X-Radar, the first extensive real-world multi-modal dataset featuring 4D Radar. In this section, we first describe the data acquisition in Sec. 3.1; then, we present the data annotation in Sec. 3.2; and finally, we delve into the diverse data distribution and dataset analysis in Sec. 3.3.
@@ 3.1 Data Acquisition
Sensor Setup. The dataset was collected using a connected vehicle-side platform (Fig. 2(a)) and an intelligent roadside unit (Fig. 2(b)). Both units are equipped with sensors, including 4D Radar, LiDAR, and multi-view cameras. Additionally, a GPS/IMU system is employed to achieve high-precision localisation, facilitating the initial point cloud registration between the vehicle-side and roadside platforms. A C-V2X unit is also integrated for wireless data transmission. The sensor layout configuration can be found in Fig. 2, while detailed specifications are listed in Tab. 2.
Figure 2: The sensor configuration on the connected vehicle-side platform and the intelligent roadside unit. a) the vehicle-side platform, and b) the intelligent roadside unit or infrastructure unit. Both are equipped with multi-modal sensors, including cameras, LiDAR, and 4D Radar, along with a C-V2X unit and a GPS/IMU system.
@@
Synchronization. For cooperative perception datasets, synchronising sensors on both vehicle-side and roadside platforms using a uniform timestamp standard is crucial. To ensure consistency, all computer clocks are initially aligned with GPS time. Hardware-triggered synchronisation of LiDAR, cameras, and 4D Radar is then implemented using the Precision Time Protocol (PTP) and Pulse Per Second (PPS) signals. Subsequently, the closest LiDAR frames from the vehicle-side and roadside platforms are matched, and the camera and 4D Radar data are aligned with each corresponding LiDAR frame to create unified multi-modal data frames. Finally, the time difference between sensors across the two platforms is kept below 20 milliseconds for each sample.
Sensor Calibration and Registration. Through the sensor calibration process, we achieve spatial synchronisation of the camera, LiDAR, and 4D Radar. The intrinsic parameters of the camera are calibrated using a checkerboard pattern. In contrast, the LiDAR is calibrated relative to the camera by utilising 100 point pairs extracted from the point cloud and the corresponding camera image. The extrinsic parameters are derived by minimising the reprojection errors between the 2D-3D point correspondences. The calibration of the LiDAR with the 4D Radar is conducted by selecting 100 high-intensity point pairs located on corner reflectors. The results of this calibration are visually represented in Fig. 3. Vehicle-side LiDAR alignment with roadside LiDAR is achieved through point cloud registration, initially computed using RTK localisation and subsequently refined via CBM [24] and manual adjustments. The visualisation of the point cloud registration is shown in Fig. 4.
Figure 3: Visualization of calibration results. a) The calibration results between the camera and LiDAR. b) The calibration results between the 4D Radar and LiDAR / Camera. The LiDAR points are projected onto the camera plane using the camera’s intrinsic parameters and the camera-LiDAR extrinsics. Similarly, the 4D Radar points are transferred to the LiDAR coordinate system using the 4D Radar-LiDAR extrinsics. Additionally, these 4D Radar points are also mapped onto the camera plane by employing the camera’s intrinsic parameters, along with the extrinsic parameters of the camera-LiDAR and 4D Radar-LiDAR.
@@
Data Collection. We collected 15 hours of driving data, comprising 540K frames and capturing various weather conditions, such as sunny, rainy, foggy, and snowy, across daytime, dusk, and nighttime, as well as typical challenging intersection scenarios. Further details can be found in the supplementary materials. From this, we manually selected the most representative 40 sequences to form V2X-Radar-C; each sequence lasts between 10 to 25 seconds, with a capturing frequency of 10 Hz. Based on this, we also sampled 10 vehicle-only sequences to create V2X-Radar-V and sampled an additional 10 infrastructure-only sequences to form V2X-Radar-I. In comparison to the single-view data in V2X-Radar-C, both V2X-Radar-V and V2X-Radar-I encompass a wider variety of scenes. Consequently, these three sub-datasets collectively included 20K LiDAR frames, 40K camera images, and 20K 4D Radar data.
@@
Coordinate System. Our dataset includes four types of coordinate systems: (1) LiDAR coordinate system, where the X, Y, and Z axes align with the front, left, and upward directions of the LiDAR. (2) Camera coordinate systems, in which the z-axis denotes depth. (3) 4D Radar coordinate system, with the X, Y, and Z axes oriented towards the right, front, and upward directions. (4) Global coordinate system, aligning with the LiDAR frame on the roadside platform.
Figure 5: Data analysis of our V2X-Radar dataset. a) Distribution of objects during day and night conditions. b) Average and maximum number of LiDAR points within the 3D bounding box for each category. c) Average and maximum number of 4D Radar points within the 3D bounding box for each category. d) Number of annotations per collaborative sample. The vertical axes of sub-plots (a)-(c) use a log scale, whilst the vertical axis of sub-plot (d) employs a standard scale.
3D Bounding Boxes Annotation. Data annotation encompasses annotations for vehicles, roadsides, and collaborative efforts. Vehicle and roadside annotations are manually created within their respective LiDAR coordinate systems. For collaborative annotation, we first align the vehicle-side and roadside annotations within a unified roadside LiDAR coordinate system, then match them using the Intersection over Union (IoU) metric to remove duplicates. The annotation process involves five object categories: pedestrian, cyclist, car, bus, and truck. Each object annotation includes attributes such as occlusion and truncation, alongside the geometric properties of 3D bounding boxes, denoted as $(x,y,z,w,h,l,\theta)$, where $(l.w.h)$ represent the object’s dimensions, $(x,y,z)$ indicate its location, and $heta$ denotes its orientation.